Usman`s father is twice as old as he is twenty years ago.Usman`s father is seven times as old as his son. Find their current ages?
Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

x = father age now
y = son age now
---
x = 2(x - 21)
x = 7y
---
x = 2x - 42
x = 42
---
x = 7y
y = x/7
y = 42/7
y = 6
---
answer:
x = father age now = 42
y = son age now = 6
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---

Answer by stanbon(75887)

Twenty years ago Usman`s father was twice as old as Usman is now
Usman`s father is now seven times as old as his son.
Find their current ages?
=====
Equations:
f-20 = 2u
f = 7*u
----
Substitute for "f" and solve for "u"::
7u-20 = 2u
5u = 20
u = 4 (son's age now)
------
f = 7u = 28 (father's age now)
--------------
Cheers,
Stan H.
===============